OpenStudy (anonymous):

25e^0.1x/e^0.1x + 3=10

OpenStudy (anonymous):

Can you type it out using the equation toolbox? This question is unclear.

OpenStudy (anonymous):

Written as it, it works out to 25+3 = 10 which makes no sense.

OpenStudy (anonymous):

yes, exactly.

OpenStudy (anonymous):

yes exactly u r true

OpenStudy (anonymous):

dhatra aditya its seem to be that u r indian

OpenStudy (anonymous):

\[25e^0.1x \div e^0.1x +3 = 10\]

OpenStudy (anonymous):

\[\frac{25e^{0.1x}}{e^{0.1x}}+3=10\]Is that what you're saying?

OpenStudy (anonymous):

Yes, sorry

OpenStudy (anonymous):

thats the same as 25 = 7.

OpenStudy (anonymous):

If what I've written is correct, there are no solutions. The statement is false.

OpenStudy (anonymous):

The 3 is written in the denominator

OpenStudy (anonymous):

AHHHH....

OpenStudy (anonymous):

But you'd have\[\frac{25}{3}=10\]

OpenStudy (anonymous):

Unless it's\[\frac{25e^{0.1x}}{e^{0.1x}+3}=10\]

OpenStudy (anonymous):

Is that what you mean ^^ ?

OpenStudy (anonymous):

yes, that

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Multiply both sides by\[e^{0.1x}+3\]to obtain\[25e^{0.1x}=10(e^{0.1x}+3)=10e^{0.1x}+30\]

OpenStudy (anonymous):

Subtract 10e^(0.1x) from both sides to get\[15e^{0.1x}=30 \rightarrow e^{0.1x}=2\]

OpenStudy (anonymous):

Now take the natural logarithm of both sides:\[\ln e^{0.1x}=\ln 2 \rightarrow 0.1x=\ln 2\]Multiply both sides by 10 to give,\[x=10\ln 2\]

OpenStudy (anonymous):

hello